OpenStudy (anonymous):

Straight line 4x-5y -40=0 is tangent to the ellipse find their meeting point (x^2)/50+(y^2)/32=1

OpenStudy (anonymous):

have u tried pemdas?

OpenStudy (anonymous):

x^2+4y=36=>(x^2)/36+(4y^2)/36=36/36=>(x^2)/36+(y^2)/9=1=>(12;3)=(x;y)=>12x/36+3y/9=1=>x+y-108=0 we should use this xx/(a^2)+yy/(b^2)=1

OpenStudy (anonymous):

good job keep going ur on the right track

OpenStudy (anonymous):

Straight line 4x-5y -40=0 is tangent to the ellipse find their meeting point (x^2)/50+(y^2)/32=1

OpenStudy (anonymous):

and then...

OpenStudy (anonymous):

The elipse equation is\[x^2/36+y^2/20=1\]find the rectangle surface MNPQ